OpenStudy (anonymous):

a − 3b = 9 a = b − 3 What is the solution to the set of equations in the form (a, b)?

OpenStudy (anonymous):

Is this slope intercept?

OpenStudy (anonymous):

or what are you looking for exactly?

OpenStudy (anonymous):

coordinates

OpenStudy (anonymous):

do you think you know the answer

OpenStudy (eric_d):

a=9+3b a=b-3

OpenStudy (texaschic101):

a - 3b = 9 a = b - 3 so sub in b - 3 in for a in the 1st equation (b - 3) - 3b = 9 -- combine like terms -2b - 3 = 9 --- add 3 to both sides -2b = 9 + 3 -2b = 12 b = 12/-2 b = -6 a = b - 3 a = -6 - 3 a = -9 check.. a - 3b = 9 -9 - 3(-6) = 9 -9 + 18 = 9 9 = 9 (correct) so a = -9 and b = -6

OpenStudy (eric_d):

substitute 1st eqn into 2nd eqn

OpenStudy (anonymous):

the choices are (−9, −6) (−4, −3) (−6, −3) (−9, −7)

OpenStudy (texaschic101):

(-9,-6)

OpenStudy (anonymous):

how did you get that

OpenStudy (texaschic101):

I did the entire problem up above ^^. Check it out and see if you have any questions ?

OpenStudy (texaschic101):

Basically, your dealing with a substitution problem

OpenStudy (anonymous):

still dont understand

OpenStudy (anonymous):

y = x + 14 y = 3x + 2

OpenStudy (texaschic101):

y = x + 14 y = 3x + 2 2 ways to do this.. (1) y = x + 14...so sub in x + 14 in for y in the other equation x + 14 = 3x + 2 x - 3x = 2 - 14 -2x = - 12 x = 6 y = x + 14 y = 6 + 14 y = 20 OR (2) y = 3x + 2....so sub in 3x + 2 in for y in the other equation 3x + 2 = x + 14 3x - x = 14 - 2 2x = 12 x = 6 y = 3x + 2 y = 3(6) + 2 y = 18 + 2 y = 20 either way you do it, the answers are the same...x = 6 and y = 20 or (6,20)

OpenStudy (texaschic101):

sorry but I have to go now...kinda lost track of time. I might be back on later...